OpenStudy (anonymous):

Use the substitution method to solve the system of equations. Choose the correct ordered pair. 4(x - 2) = 8(y - 1) 16y - 32 = 2x + 4 A. (6,3) B.(18,6) C.(6,18) D.(3,6) E.(6,1/2) Some please help ASAP please

OpenStudy (johnweldon1993):

Well first lets simplify that top equation 4(x - 2) = 8(y - 1) 4x - 8 = 8y - 8 so our 2 equations are 8y - 8 = 4x - 8 16y - 32 = 2x + 4 For the substitution method...we should solve one of these equations for 1 variable...lets choose to solve the 2nd one and we'll solve for x first 16y - 32 = 2x + 4 how would we solve this for 'x'?

OpenStudy (texaschic101):

do you need help ?

OpenStudy (texaschic101):

first you need to get the 2x on one side and everything else on the other side. Can you do that ?

OpenStudy (texaschic101):

when you decide to participate, I will help you